# Stay-Positive: A Case for Ignoring Real Image Features in Fake Image Detection

### v1

@@ Abstract
Detecting AI-generated images is a challenging yet essential task. A primary difficulty arises from the detector’s tendency to rely on spurious patterns, such as compression artifacts, which can influence its decisions. These issues often stem from specific patterns that the detector associates with the real data distribution, making it difficult to isolate the actual generative traces. We argue that an image should be classified as fake if and only if it contains artifacts introduced by the generative model. Based on this premise, we propose Stay-Positive, an algorithm designed to constrain the detector’s focus to generative artifacts while disregarding those associated with real data. Experimental results demonstrate that detectors trained with Stay-Positive exhibit reduced susceptibility to spurious correlations, leading to improved generalization and robustness to post-processing. Additionally, unlike detectors that associate artifacts with real images, those that focus purely on fake artifacts are better at detecting inpainted real images.
@@ 1 Introduction
Recent advancements in deep generative modeling, particularly diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019) and flow-based models (Liu et al., 2022a; Lipman et al., 2022), have significantly improved image generation capabilities. The combination of large-scale datasets (Schuhmann et al., 2022) and architectural innovations (Dhariwal & Nichol, 2021; Esser et al., 2024) has enabled modern systems to generate high-quality images (Pernias et al., 2023; Labs, ; Podell et al., 2023), often conditioned on auxiliary inputs such as text. While these advancements have expanded the potential applications of generative models, they also raise concerns about misuse, including misinformation and fraud. Image forensics aims to address these risks by developing methods to reliably detect AI-generated images.
Significant progress has been made in fake image detection, with detectors trained to generalize across CNN-based generators (Wang et al., 2020) and extended to unseen architectures (Ojha et al., 2023). However, reliably detecting fake images from a known generator family remains a challenge. A key difficulty arises when images undergo post-processing (e.g., compression), as demonstrated by Rajan et al. (2024), where detectors trained on images generated by latent diffusion models (Vahdat et al., 2021; Rombach et al., 2022) exhibited reduced effectiveness in detecting post-processed images. Specifically, detectors trained by Corvi et al. (2023) and Rajan et al. (2024) were found to spuriously associate WEBP compression artifacts with real images, which hurts their performance. In this work, we focus on improving the detection of fake images generated by a known model family by mitigating the impact of spurious correlations.
The most effective approach for developing a fake image detector is to train a neural network-based binary classifier. However, its performance is highly dependent on the training data, and any discriminative feature linked to it, including subtle post-processing artifacts, can spuriously influence the detector’s decisions. A common cause of such issues is the real images used for training, which are often sourced from online platforms and may have undergone unknown operations such as compression and resizing before being uploaded. As a result, detectors can learn spurious correlations. For example, models trained by Corvi et al. (2023) and Rajan et al. (2024) exhibit difficulties in handling WEBP-compressed images. Similar issues with compression artifacts were observed in the GenImage benchmark (Zhu et al., 2024), as reported by Grommelt et al. (2024).
@@ 4 Stay-Positive: Learning to Ignore Real Features
Stay-Positive to Ignore Real Features: Based on our analysis in Section 3.2, we aim for the real score to be 0 for all images. By definition, avoiding negative connections in the last layer guarantees this outcome. We achieve this by constraining the last-layer weights during optimization, overwriting any negative values with zero. This ensures that the network relies solely on $\mathcal{I}_{\text{fake}}$ to fit the training data. In our experiments, we found that training the entire network with this constraint negatively impacted the classification accuracy of real images. We hypothesize that when the entire network is trained, it may learn to link the absence of spurious real features to fake images, using negations. We discuss this further in Appendix A.4. To avoid this, we perform last-layer retraining instead. The details are outlined in Algorithm 1. The algorithm illustrates a simple SGD update, though it is independent of the optimizer used.
@@ 5 Experiments
In this section, we evaluate whether ignoring real features can improve the performance of existing detectors. We consider the following state-of-the-art baselines: The Corvi detector (Corvi et al., 2023), which is based on a ResNet-50 network trained on real images from MSCOCO and LSUN, and fake images generated by LDM with prompts corresponding to the real data. Similarly, the Rajan detector (Rajan et al., 2024) uses the same real images as Corvi, but instead trains on fake images generated by LDM reconstructions of the real images.
@@ 5.1.2 Resizing-based Artifacts
Corvi struggles with downsized fake images, as noted by Rajan et al. (2024). This issue arises from the data augmentation strategy used during training, where real images, which typically have higher resolutions, are randomly cropped and downsized to 256×256. This augmentation causes the detector to associate downsizing with real images.
@@
Table 1: AP of existing detectors before and after last-layer retraining. The large improvements on FLUX and aMUSEd suggest that existing detectors are harmed by spurious correlations unrelated to post-processing, which our algorithm circumvents. Additionally, on settings where Corvi, Rajan already perform well, our method performs the same if not slightly better.
Method SD MJ KD PG PixArt LCM FLUX Wuerstchen aMUSEd
Corvi 99.97 99.46 99.98 97.42 99.99 99.94 57.25 100 89.18
Corvidirect-$\oplus$ (Ours) 99.98 99.79 99.99 99.99 99.99 99.99 99.33 99.99 99.80
Rajan 99.89 99.90 99.98 99.94 100 99.99 80.64 95.13 87.20
Rajandirect-$\oplus$ (Ours) 99.99 99.93 99.99 99.99 99.99 99.99 90.50 97.99 98.11
@@ 5.2.1 Dataset
To test our hypothesis, we use the following dataset: i. Real: 6,000 images from Redcaps (Desai et al., 2021), WikiArt (wik, ), LAION-Aesthetics (Schuhmann et al., 2022), and whichfaceisreal (whi, ), including 3,000 post-processed images. ii. SD: 3,000 images from SDv1.5, InstructPix2Pix (Brooks et al., 2023), and Nights (Fu et al., 2023); iii. MJ: 3,000 images from MidJourney v4 and v5 (mid, ); iv. Kandinsky: 3,100 images from Kandinsky 2.1 (Razzhigaev et al., 2023); v. PG: 3,150 images from Playground v2.5 (Li et al., 2024); vi. PixArt: 3,150 images from PixelArt-$\alpha$ (Chen et al., 2023); vii. LCM: 3,146 images from SimianLuo/LCM_Dreamshaper_v7 (Luo et al., 2023); viii. FLUX: 3,000 images from Black-Forest-Labs/FLUX.1-dev (Labs, ); ix. Wuerstchen: 3,150 images from warp-ai/wuerstchen (Pernias et al., 2023); x. aMUSEd: 3,150 images from Amused/amused-512 (Patil et al., 2024).
@@ 5.2.2 Discussion
We present the AP values for the original Corvi and Rajan, alongside our versions incorporating last-layer retraining, in Table 1. Our methods show substantial improvements in AP on FLUX-generated images. Corvidirect-$\oplus$ (Ours) and Rajandirect-$\oplus$ (Ours) outperform the original detectors by 42.08 and 9.86, respectively, on FLUX, and by 10.62 and 10.91, respectively, on aMUSEd. As our fake images lack post-processing artifacts, the performance gap is likely due to the spurious correlations discussed in Section 3.2. These improvements suggest that while the original detectors can identify features distinguishing fake images from FLUX and aMUSEd, learning real features hinders the detection of these fake images. Importantly, in settings where Corvi and Rajan perform well (e.g., SD, KD, LCM), our method matches or outperforms them, showing that the detector does not lose performance by ignoring real features.
@@ 5.3 Comparison with State-of-the-Art Fake Detectors
Next, we compare the performance of our detector with state-of-the-art fake detection methods. We use the same dataset from Section 5.2, but to simulate a real-world setting, we create post-processed versions of the fake images. For FLUX, Wuerstchen, and aMUSEd, we randomly apply compression, resizing, and color jittering, and add these modified images to the dataset. Images for the other models are taken from the dataset provided by Rajan et al. (2024). Additionally, we evaluate our detector on the widely-used GenImage benchmark (Zhu et al., 2024), with results presented in Appendix A.3.
Table 2: AP of state-of-the-art detectors on post-processed fake images. Methods trained by our algorithm, Corvidirect-$\oplus$ and Rajandirect-$\oplus$ outperform their base detectors as well as other state-of-the-art detection approaches on recent generators such as FLUX and Wuerstchen. Additionally, on settings where Corvi, Rajan already perform well, our method performs the same if not slightly better.
Method SD MJ KD PG PixArt LCM FLUX Wuerstchen aMUSEd
UFD-ProGAN (Ojha et al., 2023) 61.93 61.72 74.88 70.23 69.81 70.86 37.54 86.30 88.84
UFD-LDM (Ojha et al., 2023) 62.02 52.33 65.33 62.36 62.39 65.52 35.18 86.46 90.98
ClipDet (Cozzolino et al., 2024) 71.63 73.72 74.71 75.53 76.61 72.06 81.48 90.11 87.61
DRCT (Chen et al., 2024) 96.09 90.74 95.86 83.83 78.39 88.43 76.60 85.40 87.87
Corvi (Corvi et al., 2023) 97.87 94.81 95.38 91.40 94.45 96.33 74.57 95.83 94.57
Corvidirect-$\oplus$ (Ours) 98.94 94.92 97.71 97.87 98.59 98.73 94.23 98.16 95.47
Rajan (Rajan et al., 2024) 99.40 98.30 98.18 98.62 98.64 99.79 87.80 94.51 95.38
Rajandirect-$\oplus$ (Ours) 99.22 96.98 98.22 98.53 99.11 99.57 91.85 94.74 97.26
@@ 5.3.1 Baselines
We select 3 methods from the CLIP linear-probing paradigm: i. UFD-ProGAN (Ojha et al., 2023): trained on ProGAN images, representing the baseline method in this paradigm; ii. UFD-LDM (Ojha et al., 2023): trained on LDM images, extending the original setup to a different generative domain; iii. ClipDet: a follow-up work by Cozzolino et al. (2024) that refines the linear-probing approach for LDM images. Additionally, we include iv. DRCT: a recent method by Chen et al. (2024) that employs DDIM inversion to reconstruct both real and fake images, combined with a contrastive training objective for better generalization, using a ConvNext (Liu et al., 2022b) backbone trained on images from SDv1.4. We also compare against the original Corvi and Rajan, the detectors on which our method is based, to benchmark the improvements our approach introduces.
@@ 5.3.2 Discussion
We present our results in Table 2 and observe that detectors relying on full-network fine-tuning, such as DRCT, Corvi, and Rajan, along with our improved baselines, outperform CLIP-based approaches. In comparison to Corvi, our Corvidirect-$\oplus$ (Ours) shows large improvements. On PG and FLUX generated images, our version outperforms the original Corvi by 6.47 and 19.66 respectively. On FLUX generated images, our version outperforms the original Rajan by 4.05. Notably, the AP scores of the original Corvi and Rajan on FLUX and aMUSEd improve compared to the values reported in Table 1, indicating that post-processed fake images from these generators were actually easier to detect compared to fake images without post-processing. This suggests that the fake images used in Table 1 contain artifacts that harm the performance of Corvi and Rajan. Post-processing these images can remove these artifacts, improving both detectors’ performance. Given their sensitivity to post-processing, these are likely low-level artifacts affecting detector performance, similar to those described in Section.3.2. Unlike the original Corvi and Rajan, our versions are unaffected by such discrepancies.
@@ 5.4.2 Analysis
We report our results in Table 3. At the 50-level, all approaches barring Corvidirect-$\oplus$ and Rajandirect-$\oplus$ fail to detect these inpainted images. This is perhaps unsurprising since these images are ”real” for the most part. Our detectors on the other hand, do not suffer from these issues as indicated by the AP score. The performance of Corvi and Rajan improves with more inpainting, highlighting the harmful impact of real features on their performance.

### v2

@@ Abstract
Detecting AI-generated images is a challenging yet essential task. A primary difficulty arises from the detector’s tendency to rely on spurious patterns, such as compression artifacts, which can influence its decisions. These issues often stem from specific patterns that the detector associates with the real data distribution, making it difficult to isolate the actual generative traces. We argue that an image should be classified as fake if and only if it contains artifacts introduced by the generative model. Based on this premise, we propose Stay-Positive, an algorithm designed to constrain the detector’s focus to generative artifacts while disregarding those associated with real data. Experimental results demonstrate that detectors trained with Stay-Positive exhibit reduced susceptibility to spurious correlations, leading to improved generalization and robustness to post-processing. Additionally, unlike detectors that associate artifacts with real images, those that focus purely on fake artifacts are better at detecting inpainted real images. For implementation details, please visit: https://anisundar18.github.io/Stay-Positive.
@@ 1 Introduction
Recent advancements in deep generative modeling, particularly diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019) and flow-based models (Liu et al., 2023; Lipman et al., 2022), have significantly improved image generation capabilities. The combination of large-scale datasets (Schuhmann et al., 2022) and architectural innovations (Dhariwal & Nichol, 2021; Esser et al., 2024) has enabled modern systems to generate high-quality images (Pernias et al., 2023; Labs, ; Podell et al., 2024), often conditioned on auxiliary inputs such as text. While these advancements have expanded the potential applications of generative models, they also raise concerns about misuse, including misinformation and fraud. Image forensics aims to address these risks by developing methods to reliably detect AI-generated images.
Significant progress has been made in fake image detection, with detectors trained to generalize across CNN-based generators (Wang et al., 2020) and extended to unseen architectures (Ojha et al., 2023). However, reliably detecting fake images from a known generator family remains a challenge. A key difficulty arises when images undergo post-processing (e.g., compression), as demonstrated by Rajan et al. (2025), where detectors trained on images generated by latent diffusion models (Vahdat et al., 2021; Rombach et al., 2022) exhibited reduced effectiveness in detecting post-processed images. Specifically, detectors trained by Corvi et al. (2023) and Rajan et al. (2025) were found to spuriously associate WEBP compression artifacts with real images, which negatively impacted their performance. In this work, we focus on improving the detection of fake images generated by a known model family by mitigating the impact of spurious correlations. The most effective approach for developing a fake image detector is to train a neural network-based binary classifier. However, its performance is highly dependent on the training data, and any discriminative feature associated with the data, including subtle post-processing artifacts, can influence the detector’s decisions. A common source of such issues is the use of real images for training, which are often collected from online platforms and may have undergone unknown operations such as compression and resizing prior to upload. As a result, detectors can learn spurious correlations. Such issues with compression artifacts were also observed in the widely adopted GenImage benchmark (Zhu et al., 2024), as reported by Grommelt et al. (2024).
@@ 4 Stay-Positive: Learning to Ignore Real Features
Stay-Positive to Ignore Real Features: Based on our analysis in Section 3.2, we aim for the real score to be 0 for all images. By definition, avoiding negative connections in the last layer guarantees this outcome. We achieve this by constraining the last-layer weights during optimization, overwriting any negative values with zero. This ensures that the network relies solely on $\mathcal{I}_{\text{fake}}$ to fit the training data. In our experiments, we found that training the entire network with this constraint negatively impacted the classification accuracy of real images. We report these results in Section 5.5. We hypothesize that when the entire network is trained, it may learn to link the absence of spurious real features to fake images, using negations. We discuss this further in Appendix A.6. To avoid this, we perform last-layer retraining instead. The details are outlined in Algorithm 1. The algorithm illustrates a simple SGD update, though it is independent of the optimizer used.
@@ 5 Experiments
In this section, we evaluate whether ignoring real features can improve the performance of existing detectors. We consider the following state-of-the-art baselines: The Corvi detector (Corvi et al., 2023), which is based on a ResNet-50 network trained on real images from MSCOCO and LSUN, and fake images generated by LDM with prompts corresponding to the real data. Similarly, the Rajan detector (Rajan et al., 2025) uses the same real images as Corvi, but instead trains on fake images generated by LDM reconstructions of the real images.
@@ 5.1.2 Resizing-based Artifacts
Corvi struggles with downsized fake images, as noted by Rajan et al. (2025). This issue arises from the data augmentation strategy used during training, where real images, which typically have higher resolutions, are randomly cropped and downsized to 256×256. This augmentation causes the detector to associate downsizing with real images. We also repeat this experiment using images from the Synthbuster dataset (Cozzolino et al., 2024), details of which can be found in Appendix A.3.
@@
Table 1: AP of existing detectors before and after last-layer retraining. The large improvements on FLUX and aMUSEd suggest that existing detectors are harmed by spurious correlations unrelated to post-processing, which our algorithm circumvents. Additionally, on settings where Corvi, Rajan already perform well, our method performs the same if not slightly better. AVG is the equally weighted AP average across all generators.
Method SD MJ KD PG PixArt LCM FLUX Wuerstchen aMUSEd AVG
Corvi 99.97 99.46 99.98 97.42 99.99 99.94 57.25 100 89.18 94.23
Corvidirect-$\oplus$ (Ours) 99.98 99.79 99.99 99.99 99.99 99.99 99.33 99.99 99.80 99.88
Rajan 99.89 99.90 99.98 99.94 100 99.99 80.64 95.13 87.20 96.22
Rajandirect-$\oplus$ (Ours) 99.99 99.93 99.99 99.99 99.99 99.99 90.50 97.99 98.11 98.65
@@ 5.2.1 Dataset
To test our hypothesis, we use the following dataset: i. Real: 6,000 images from Redcaps (Desai et al., 2021), WikiArt (wik, ), LAION-Aesthetics (schuhmann2022laion), and whichfaceisreal (whi, ), including 3,000 post-processed images. ii. SD: 3,000 images from SDv1.5, InstructPix2Pix (Brooks et al., 2023), and Nights (Fu et al., 2023); iii. MJ: 3,000 images from MidJourney v4 and v5 (mid, ); iv. Kandinsky: 3,100 images from Kandinsky 2.1 (Razzhigaev et al., 2023); v. PG: 3,150 images from Playground v2.5 (Li et al., 2024); vi. PixArt: 3,150 images from PixelArt-$\alpha$ (Chen et al., 2024b); vii. LCM: 3,146 images from SimianLuo/LCM_Dreamshaper_v7 (Luo et al., 2023); viii. FLUX: 3,000 images from Black-Forest-Labs/FLUX.1-dev (Labs, ); ix. Wuerstchen: 3,150 images from warp-ai/wuerstchen (Pernias et al., 2023); x. aMUSEd: 3,150 images from Amused/amused-512 (Patil et al., 2024).
@@ 5.2.2 Discussion
We present the AP values for the original Corvi and Rajan, alongside our versions incorporating last-layer retraining, in Table 1. Our methods show substantial improvements in AP on FLUX-generated images. Corvidirect-$\oplus$ (Ours) and Rajandirect-$\oplus$ (Ours) outperform the original detectors by 42.08 and 9.86, respectively, on FLUX, and by 10.62 and 10.91, respectively, on aMUSEd. As our fake images lack post-processing artifacts, the performance gap is likely due to the spurious correlations discussed in Section 3.2. These improvements suggest that while the original detectors can identify features distinguishing fake images from FLUX and aMUSEd, learning real features hinders the detection of these fake images. Importantly, in settings where Corvi and Rajan perform well (e.g., SD, KD, LCM), our method matches or outperforms them, showing that the detector does not lose performance by ignoring real features. The rightmost column reports the mean average precision aggregated over all settings.
@@ 5.3 Comparison with State-of-the-Art Fake Detectors
Next, we compare the performance of our detector with state-of-the-art fake detection methods. We use the same dataset from Section 5.2, but to simulate a real-world setting, we create post-processed versions of the fake images. For FLUX, Wuerstchen, and aMUSEd, we randomly apply compression, resizing, and color jittering, and add these modified images to the dataset. Images for the other models are taken from the dataset provided by Rajan et al. (2025). Additionally, we evaluate our detector on the widely used GenImage benchmark (Zhu et al., 2024), with results presented in Appendix A.4, as well as on latent diffusion and autoregressive models from the UFD benchmark (Ojha et al., 2023), with results shown in Appendix A.5.
Table 2: AP of state-of-the-art detectors on post-processed fake images. Methods trained by our algorithm, Corvidirect-$\oplus$ and Rajandirect-$\oplus$ outperform their base detectors as well as other state-of-the-art detection approaches on recent generators such as FLUX and Wuerstchen. Additionally, on settings where Corvi, Rajan already perform well, our method performs the same if not slightly better. AVG is the equally weighted AP average across all generators.
Method SD MJ KD PG PixArt LCM FLUX Wuerstchen aMUSEd AVG
AEROBLADE (Ricker et al., 2024) 90.81 96.48 94.03 71.53 87.84 89.99 60.34 85.93 88.39 85.03
UFD-ProGAN (Ojha et al., 2023) 61.93 61.72 74.88 70.23 69.81 70.86 37.54 86.30 88.84 64.73
UFD-LDM (Ojha et al., 2023) 62.02 52.33 65.33 62.36 62.39 65.52 35.18 86.46 90.98 69.12
ClipDet (Cozzolino et al., 2024) 71.63 73.72 74.71 75.53 76.61 72.06 81.48 90.11 87.61 78.16
DRCT (Chen et al., 2024a) 96.09 90.74 95.86 83.83 78.39 88.43 76.60 85.40 87.87 87.02
Corvi (Corvi et al., 2023) 97.87 94.81 95.38 91.40 94.45 96.33 74.57 95.83 94.57 92.80
Corvidirect-$\oplus$ (Ours) 98.94 94.92 97.71 97.87 98.59 98.73 94.23 98.16 95.47 97.17
Rajan (Rajan et al., 2025) 99.40 98.30 98.18 98.62 98.64 99.79 87.80 94.51 95.38 96.73
Rajandirect-$\oplus$ (Ours) 99.22 96.98 98.22 98.53 99.11 99.57 91.85 94.74 97.26 96.96
@@ 5.3.1 Baselines
From the family of zero-shot fake image detectors we select i. AEROBLADE (Ricker et al., 2024). From the CLIP linear-probing paradigm, we select: ii. UFD-ProGAN (Ojha et al., 2023): trained on ProGAN images, representing the baseline method in this paradigm; iii. UFD-LDM (Ojha et al., 2023): trained on LDM images, extending the original setup to a different generative domain; iv. ClipDet: a follow-up work by Cozzolino et al. (2024) that refines the linear-probing approach for LDM images. Additionally, we include v. DRCT: a recent method by Chen et al. (2024a) that employs DDIM inversion to reconstruct both real and fake images, combined with a contrastive training objective for better generalization, using a ConvNext (Liu et al., 2022) backbone trained on images from SDv1.4. We also compare against the original Corvi and Rajan, the detectors on which our method is based, to benchmark the improvements our approach introduces.
@@ 5.3.2 Discussion
We present our results in Table 2 and observe that detectors relying on full-network fine-tuning, such as DRCT, Corvi, and Rajan, along with our improved baselines, outperform CLIP-based and zero-shot approaches. In comparison to Corvi, our Corvidirect-$\oplus$ (Ours) shows large improvements. On PG and FLUX generated images, our version outperforms the original Corvi by 6.47 and 19.66 respectively. On FLUX generated images, our version outperforms the original Rajan by 4.05. Notably, the AP scores of the original Corvi and Rajan on FLUX and aMUSEd improve compared to the values reported in Table 1, indicating that post-processed fake images from these generators were actually easier to detect compared to fake images without post-processing. This suggests that the fake images used in Table 1 contain artifacts that harm the performance of Corvi and Rajan. Post-processing these images can remove these artifacts, improving both detectors’ performance. Given their sensitivity to post-processing, these are likely low-level artifacts affecting detector performance, similar to those described in Section.3.2. Unlike the original Corvi and Rajan, our versions are unaffected by such discrepancies.
@@ 5.4.2 Analysis
We report our results in Table 3. At the 50-level, all approaches barring Corvidirect-$\oplus$ and Rajandirect-$\oplus$ fail to detect these inpainted images. This is perhaps unsurprising since these images are “real” for the most part. Our detectors on the other hand, do not suffer from these issues as indicated by the AP score. The performance of Corvi and Rajan improves with more inpainting, highlighting the harmful impact of real features on their performance.
5.5 Ablations
In this section, we ablate other possible alternatives to final-layer re-training. We experimented with two other variants, (i) + clamped (no retrain) where we clamp the detector’s weights to stay-positive without re-training the whole network and (ii) + clamped (retrain): where we re-train the whole network with the stay-positive algorithm applied on the final layer. We calculate the average precision (AP) and report the values in Table 4.
Table 4: AP for different ablations. We experiment with different ways of applying the stay-positive algorithm. We observe that in both settings, freezing the backbone of the network while re-training the last-layer performs the best. AVG is the equally weighted AP average across all generators.
Method SD MJ KD PG PixArt LCM FLUX Wuerstchen aMUSEd AVG
Corvi 97.87 94.81 95.38 91.40 94.46 96.33 74.57 95.83 94.57 92.80
Corvi + clamped (no retrain) 85.75 76.44 85.41 82.28 93.21 85.12 63.60 85.71 66.70 80.47
Corvi + clamped (retrain) 98.93 95.38 95.42 97.64 94.98 95.54 69.50 96.76 95.80 93.33
Corvidirect-$\oplus$ (Ours) 98.94 94.92 97.71 97.87 98.59 98.73 94.23 98.16 95.47 97.18
Rajan 99.40 98.30 98.18 98.62 98.64 99.79 87.80 94.51 95.38 96.73
Rajan + clamped (no retrain) 95.01 90.83 92.72 94.57 98.87 96.50 70.47 81.71 72.85 88.17
Rajan + clamped (retrain) 99.19 96.30 95.49 98.66 95.93 96.95 58.68 88.73 94.66 91.62
Rajandirect-$\oplus$ (Ours) 99.22 96.98 98.22 98.53 99.11 99.57 91.85 94.74 97.26 97.28
We evaluate these ablated models under the same setting described in Section 5.3. Our results, similar to those in Table 2, show that clamping without retraining leads to suboptimal performance, likely due to improper reweighting of fake features. Training the entire backbone while clamping the final layer underperforms on FLUX images, likely due to the emergence of newly learned spurious fake features as discussed in Appendix A.6.
5.6 Robustness to other Post-Processing Artifacts
We also conduct tests testing the senstitivty of our detector to post-processing operations. We use JPEG Compression, additive gaussian noise and low-pass filtering to account for some of the common post-processing operations. We report our results in Appendix A.3.